# Black-Box On-Policy Distillation of Large Language Models

### v2

@@ Abstract
Figure 1: Comparison between GAD and sequence-level knowledge distillation (SeqKD; skd) trained on LMSYS-Chat [lmsys] dataset, evaluated by averaged GPT-4o scores. Left: Results on the LMSYS-Chat test set. Right: Average performance across Dolly [dolly], SelfInst [self_inst], and Vicuna [vicuna] datasets.
@@ 1 Introduction
Knowledge distillation [kd] in large language models (LLMs; gpt4, gpt5, deepseekv3, qwen3) is primarily used to create smaller, more efficient student models that retain much of the performance of a larger, resource-intensive teacher model. The setting in which the student has access to the teacher’s internal probability distribution or hidden states is called white-box distillation. Standard white-box approaches align the teacher and student by matching their output distributions, typically via Kullback-Leibler divergence (KLD) [lightpaff, minillm], or their inner states [tinybert, bert-pkd, minilm]. However, white-box access is often impractical when the teacher is a proprietary API model (e.g., GPT-5). In this scenario, only teacher-generated texts are accessible, defining the more challenging black-box distillation setting. The absence of fine-grained probability supervision makes conventional likelihood-based objectives unavailable. Typical black-box distillation methods simply perform supervised fine-tuning on teacher responses [alpaca, vicuna]. Furthermore, when the student and teacher employ incompatible tokenizers, applying likelihood-based objectives also becomes challenging. This highlights the need for a framework that can effectively extract deeper and richer knowledge from teacher-generated text responses.
@@
Recent studies [minillm, googlepolicy, thinkingmachine-onpolicy, qwen3] in white-box distillation highlight the importance of on-policy learning, where the student learns from its own generated responses rather than solely imitating the teacher’s outputs. These studies show that performing reverse KLD on student-generated text promotes mode-seeking behavior and reduces exposure bias compared to teacher-forced training. However, extending this idea to the black-box setting introduces a major challenge: when the student produces its own responses, there are no probability-level supervision signals available from the teacher to evaluate or correct them. Without explicit feedback, the student cannot directly gauge the quality of its generations relative to the teacher, making effective on-policy distillation infeasible under the standard likelihood-based framework.
@@
To address this limitation, we propose GAD, a Generative Adversarial Distillation framework that enables on-policy learning in the black-box regime. Our key idea is to view the student as a generator that produces responses conditioned on prompts, and to train a discriminator to distinguish between teacher and student outputs. The generator is then optimized to produce responses that the discriminator cannot distinguish from those of the teacher, forming a minimax game similar to generative adversarial networks (GANs; gan, seqgan). This adversarial process allows the student to receive implicit feedback on the quality of its own generations, even without access to the teacher’s probability space. Besides, from the perspective of reinforcement learning (RL; rlintro, ppo, trpo), our discriminator can be interpreted as an on-policy reward model that evolves jointly with the student policy. Unlike conventional reward models in RLHF [instruct-gpt] which are fixed after pretraining and prone to reward hacking [reward_hacking], our discriminator continually adapts to the student’s behavior during training. The on-policy reward modeling provides stable and dynamic supervision throughout the distillation process.
@@
We validate our approach using GPT-5-Chat [gpt5] as a teacher and a range of open-source models from the Qwen2.5 [qwen2.5] and Llama3 [llama3] families as a student. Experiments are conducted on the a subset of LMSYS-Chat-1M dataset [lmsys] and evaluated across multiple domains. Under identical training budgets, GAD consistently outperforms both the instruction models before distillation and the SeqKD [skd, vicuna, alpaca, ITGPT4, lima] baseline across all datasets and model sizes. Notably, on GPT-4o score, Qwen2.5-3B-Instruct distilled with GAD matches the performance of Qwen2.5-7B-Instruct distilled with SeqKD, while Qwen2.5-14B-Instruct trained with GAD approaches the capability of the GPT-5 teacher itself. Our method also delivers particularly strong improvements in out-of-distribution generalization, where SeqKD yields marginal or negative gains. Human evaluations further confirm performance. GAD can effectively extract high-quality knowledge from black-box LLMs without access to output logits.
@@ 2 Method
We study conditional text generation of large language models, where a model generates a response $y$ conditioned on a given prompt $x$ sampled from dataset $\mathcal{T}$. To transfer the capabilities of large models to smaller ones, knowledge distillation (KD) trains a student distribution $q_{\theta}(y|x)$ parameterized by $\theta$ to approximate the behavior of a teacher distribution $p(y|x)$. In the white-box distillation setting, the student has access to the teacher’s predictive distribution $p(y|x)$. Approaches such as forward KLD [skd, lightpaff, vicuna, alpaca] or reverse KLD [minillm] are designed for this setting. However, these techniques can fail if the teacher is a proprietary model that only returns generated text. We refer to this scenario as black-box distillation, where only textual responses from the teacher are observable. The goal is to learn a student model that imitates the teacher’s generative behavior without access to its internal probability space.
@@ 2.1 GAD: Generative Adversarial Distillation
We perform black-box distillation with generative adversarial training [gan, seqgan] as shown in Figure˜2. The training dataset $\mathcal{T}=\{(x,y_{t})\}$ is constructed by iterating over the prompts $x$ in the original dataset and sampling a teacher response $y_{t}$ for each. Our framework consists of a generator $G$ which is the student model, and a discriminator $D$ that assesses the quality of the student and teacher responses. The generator generates the response $G(x)$ to the prompt $x$. The discriminator predicts a sequence-level scalar score $D([x,y])$ given prompt $x$ and response yy111 The input prompt $x$ and generated response $y$ are concatenated (i.e., $[x,y]$) and fed into the discriminator (i.e., $D([x,y])$). For brevity, we use $D(y)$ below to represent $D([x,y])$.. The discriminator is initialized using generator model parameters with an extra prediction head. The head projects the final hidden state to a scalar score, and the score of the last token in the sequence is taken as the sequence-level score. The training objective is formulated as a two-player minimax game with the following value function $\mathcal{V}(G,D)$:
@@
where $\sigma(\cdot)$ denotes the sigmoid function. We use Bradley-Terry model [bradley] to capture pairwise preferences between teacher and student response. The proposed generative adversarial training framework allows the student to learn on-policy from its own generated responses via discriminator feedback, eliminating the need to access the teacher’s internal representations.
@@ 2.2 Training
Since the sampling operation in $G(x)$ is non-differentiable with respect to the student model parameters, we treat $D(G(x))$ as a reward and optimize it using policy gradient [policy_gradient] with established reinforcement learning algorithms. We employ GRPO [grpo] to train the student in our experiments, with detailed formulations provided in Appendix A.1. For the discriminator $D$, we minimize its training loss derived from Equation (1):
@@ 2.3 Implement GAD with Reinforcement Learning Frameworks
In our experiments, we implement GAD using existing reinforcement learning frameworks, such as verl [verl]. GRPO [grpo] is used as the policy gradient algorithm, which is detailed in Appendix A.1.
@@
Unlike vanilla reinforcement learning, GAD also needs to jointly update the discriminator (i.e., reward model). The discriminator is trained with Bradley-Terry loss on preference pairs to score the teacher response higher than the student’s output, similar to the reward model in RLHF [instruct-gpt]. While conventional RLHF trains a fixed reward model prior to policy optimization which is prone to reward hacking, our approach updates the reward model (discriminator) online to adapt it to the current policy continually.
@@
Term
Correspondence
Policy Model Generator (i.e.,Student LLM)
@@ 3.1 Setup
Given a dataset of instruction prompts, we collect corresponding responses from a teacher model and use them to distill student models. For the following experiments, we use LMSYS-Chat-1M-Clean222 https://huggingface.co/datasets/OpenLeecher/lmsys_chat_1m_clean, a clean version of the LMSYS-Chat-1M dataset [lmsys]. The dataset is derived from high-quality conversational data collected via the Chatbot Arena333 https://lmarena.ai platform.
@@
We adopt GPT-5-Chat [gpt5] as the teacher model. It is a closed-source chat model ranked ninth on the Chatbot Text Arena leaderboard at the time of writing. For student models, we use the instruction-tuned variants of open-source models from the Qwen2.5 [qwen2.5] family (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct) and the Llama3 [llama3] family (Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct).
@@
For training data, we sample 200K samples from LMSYS-Chat-1M-Clean and collect the corresponding GPT-5-Chat responses to the instructions as teacher responses. All models are trained for 3 epochs with a batch size of 256, totaling approximately 2400 optimization steps. The PPO mini-batch size for each policy update is also 256. The maximum context length is set to 2048 tokens for instruction prompts and 1536 for model responses. The training and sampling temperature is set to $0.8$. We save checkpoints every 50 steps. More training details can be found in Appendix A.2.
@@
We reserve 500 samples of LMSYS-Chat-1M-Clean as the primary test set. We also include test datasets consisting of a 500-sample subset split from Dolly [dolly], the 252-sample SelfInst dataset [self_inst], and the 80-question Vicuna benchmark [vicuna] to evaluate out-of-distribution generalization. We report the GPT-4o evaluation scores [mtbench, minillm], where GPT-4o first generates reference answers and then scores the output of the student model against them. We also conduct human evaluations on the LMSYS-Chat-1M-Clean test set for qualitative assessment. We select the checkpoint that achieved the highest GPT-4o score and whose response length is within an acceptable range for each experiment. Detailed evaluation protocols are described in Appendix A.3.
@@ 3.2 Main Results
We report the results of automatic evaluation using GPT-4o scores in Figure˜1 and Table˜2. We compare GAD with the instruct model before distillation and the SeqKD baseline. Across all datasets, GAD consistently outperforms the baselines. As shown in Figure˜1, on the LMSYS-Chat test set, Qwen2.5-3B-Instruct trained with GAD matches the performance of Qwen2.5-7B-Instruct trained with SeqKD; similarly, Qwen2.5-7B-Instruct with GAD rivals Qwen2.5-14B-Instruct with SeqKD, and Qwen2.5-14B-Instruct with GAD is comparable to the GPT-5-Chat teacher. In addition, GAD shows particularly strong gains on out-of-distribution generalization benchmarks. On Dolly, SelfInst, and Vicuna, SeqKD yields marginal or even negative improvements, whereas GAD maintains robust performance gains. We attribute this to the superior generalization ability of reinforcement learning compared to supervised fine-tuning [sftmemrlgen, dft]. We also provide additional automatic evaluation results in Section˜B.1.
@@ 3.3 Analysis
We evaluate the similarity of local patterns between the student and teacher on the LMSYS-Chat test set in Figure˜5, measured by the F1 score of N-gram overlap. The student is trained from Qwen2.5-14B-Instruct, and the teacher is GPT-5-Chat. The SeqKD student exhibits a higher N-gram overlap while a lower GPT-4o evaluation score compared to the GAD student. This suggests that supervised fine-tuning tends to memorize local lexical patterns [sftmemrlgen, dft], whereas our RL-based approach better captures the teacher’s global stylistic characteristics.
@@
We simulate the optimizing patterns of GAD and SeqKD in a toy experiment shown in Figure˜5. We observe that GAD tends to learn reachable modes of the teacher, whereas SeqKD aims to cover all modes. The setup simulates a black-box distillation scenario. We define a discrete Gaussian mixture distribution as a teacher distribution $p$, which has categorical outputs $0,\dots,9$}. A student, modeled as a single Gaussian distribution, learns to imitate the teacher using only output samples without access to $p$. We compare two student training schemes, SeqKD and GAD. The GAD student is optimized using the REINFORCE algorithm [reinforce]. As illustrated in Figure˜5, the SeqKD student exhibits a mode-covering behavior, spreading probability mass across all possible outputs [minillm]. In contrast, the GAD student focuses on mode-seeking, concentrating probability optimization on reachable regions. We find that such mode-seeking behavior leads to more effective knowledge distillation in LLMs.
@@ 3.4 Ablations
We ablate the choice of discriminator loss in Section˜3.4, and observe that our default Bradley-Terry loss outperforms cross-entropy loss in overall GPT-4o evaluation score. The Bradley-Terry loss is defined in Equation (3), while the cross-entropy loss is a binary classification loss commonly adopted for discriminator in prior works [gan, seqgan, GAIL, learningdense]. The cross-entropy discriminator loss can be written as:
@@ 4 Related Work
White-box knowledge distillation of LLM assumes full access to the internal representations or token-level probabilities of a teacher model. Standard white-box approaches align the forward KLD of distribution [mixkd, lightpaff], reverse KLD of distribution [minillm], hidden states [tinybert, bert-pkd] or attention scores [minilm, minilmv2] between the teacher and the student. Recent work [minillm, thinkingmachine-onpolicy, googlepolicy] also proves the importance of on-policy distillation where the student learns from its own responses. Such approaches effectively compress large models while preserving semantic similarity. Despite their effectiveness, these methods rely on full teacher access, which is impractical for proprietary LLMs and limits their applicability to closed-source or API-only teachers.
@@
Black-box distillation trains a student model using only the textual outputs of a teacher, typically obtained by API queries to closed-source models such as GPT-5 and Gemini 2.5 [gpt5, gemini2d5]. In this setting, conventional white-box distillation methods become infeasible because of the lack of access to the teacher’s logits or hidden representations. The standard approach for this scenario, SeqKD, performs supervised fine-tuning (SFT) on the teacher’s responses [skd, ITGPT4, lima, alpaca, vicuna] to imitate the teacher’s behaviors. Recent work [s1, openthoughts, limo, deepseekr1] extends this paradigm by performing SFT on the teacher’s reasoning traces to improve the student’s reasoning ability.

### v3

@@ Abstract
Figure 1: Comparison between GAD and sequence-level knowledge distillation (SeqKD; 16) trained on LMSYS-Chat [44] dataset, evaluated by averaged GPT-4o scores. Left: Results on the LMSYS-Chat test set. Right: Average performance across Dolly [6], SelfInst [37], and Vicuna [3] datasets.
@@ 1 Introduction
Knowledge distillation [13] in large language models (LLMs; 21, 22, 18, 40) is primarily used to create smaller, more efficient student models that retain much of the performance of a larger, resource-intensive teacher model. The setting in which the student has access to the teacher’s internal probability distribution or hidden states is called white-box distillation. Standard white-box approaches align the teacher and student by matching their output distributions, typically via Kullback-Leibler divergence (KLD) [30, 10], or their inner states [15, 31, 36]. However, white-box access is often impractical when the teacher is a proprietary API model (e.g., GPT-5). In this scenario, only teacher-generated texts are accessible, defining the more challenging black-box distillation setting. The absence of fine-grained probability supervision makes conventional likelihood-based objectives unavailable. Typical black-box distillation methods simply perform supervised fine-tuning on teacher responses [34, 3]. Furthermore, when the student and teacher employ incompatible tokenizers, applying likelihood-based objectives also becomes challenging. This highlights the need for a framework that can effectively extract deeper and richer knowledge from teacher-generated text responses.
@@
Recent studies [10, 1, 19, 40] in white-box distillation highlight the importance of on-policy learning, where the student learns from its own generated responses rather than solely imitating the teacher’s outputs. These studies show that performing reverse KLD on student-generated text promotes mode-seeking behavior and reduces exposure bias compared to teacher-forced training. However, extending this idea to the black-box setting introduces a major challenge: when the student produces its own responses, there are no probability-level supervision signals available from the teacher to evaluate or correct them. Without explicit feedback, the student cannot directly gauge the quality of its generations relative to the teacher, making effective on-policy distillation infeasible under the standard likelihood-based framework.
@@
To address this limitation, we propose GAD, a Generative Adversarial Distillation framework that enables on-policy learning in the black-box regime. Our key idea is to view the student as a generator that produces responses conditioned on prompts, and to train a discriminator to distinguish between teacher and student outputs. The generator is then optimized to produce responses that the discriminator cannot distinguish from those of the teacher, forming a minimax game similar to generative adversarial networks (GANs; 8, 43). This adversarial process allows the student to receive implicit feedback on the quality of its own generations, even without access to the teacher’s probability space. Besides, from the perspective of reinforcement learning (RL; 32, 26, 25), our discriminator can be interpreted as an on-policy reward model that evolves jointly with the student policy. Unlike conventional reward models in RLHF [23] which are fixed after pretraining and prone to reward hacking [29], our discriminator continually adapts to the student’s behavior during training. The on-policy reward modeling provides stable and dynamic supervision throughout the distillation process.
@@
We validate our approach using GPT-5-Chat [22] as a teacher and a range of open-source models from the Qwen2.5 [41] and Llama3 [9] families as a student. Experiments are conducted on the a subset of LMSYS-Chat-1M dataset [44] and evaluated across multiple domains. Under identical training budgets, GAD consistently outperforms both the instruction models before distillation and the SeqKD [16, 3, 34, 24, 46] baseline across all datasets and model sizes. Notably, on GPT-4o score, Qwen2.5-3B-Instruct distilled with GAD matches the performance of Qwen2.5-7B-Instruct distilled with SeqKD, while Qwen2.5-14B-Instruct trained with GAD approaches the capability of the GPT-5 teacher itself. Our method also delivers particularly strong improvements in out-of-distribution generalization, where SeqKD yields marginal or negative gains. Human evaluations further confirm performance. GAD can effectively extract high-quality knowledge from black-box LLMs without access to output logits.
@@ 2 Method
We study conditional text generation of large language models, where a model generates a response $y$ conditioned on a given prompt $x$ sampled from dataset $\mathcal{T}$. To transfer the capabilities of large models to smaller ones, knowledge distillation (KD) trains a student distribution $q_{\theta}(y|x)$ parameterized by $\theta$ to approximate the behavior of a teacher distribution $p(y|x)$. In the white-box distillation setting, the student has access to the teacher’s predictive distribution $p(y|x)$. Approaches such as forward KLD [16, 30, 3, 34] or reverse KLD [10] are designed for this setting. However, these techniques can fail if the teacher is a proprietary model that only returns generated text. We refer to this scenario as black-box distillation, where only textual responses from the teacher are observable. The goal is to learn a student model that imitates the teacher’s generative behavior without access to its internal probability space.
@@ 2.1 GAD: Generative Adversarial Distillation
We perform black-box distillation with generative adversarial training [8, 43] as shown in Figure˜2. The training dataset $\mathcal{T}=\{(x,y_{t})\}$ is constructed by iterating over the prompts $x$ in the original dataset and sampling a teacher response $y_{t}$ for each. Our framework consists of a generator $G$ which is the student model, and a discriminator $D$ that assesses the quality of the student and teacher responses. The generator generates the response $G(x)$ to the prompt $x$. The discriminator predicts a sequence-level scalar score $D([x,y])$ given prompt $x$ and response yy111 The input prompt $x$ and generated response $y$ are concatenated (i.e., $[x,y]$) and fed into the discriminator (i.e., $D([x,y])$). For brevity, we use $D(y)$ below to represent $D([x,y])$.. The discriminator is initialized using generator model parameters with an extra prediction head. The head projects the final hidden state to a scalar score, and the score of the last token in the sequence is taken as the sequence-level score. The training objective is formulated as a two-player minimax game with the following value function $\mathcal{V}(G,D)$:
@@
where $\sigma(\cdot)$ denotes the sigmoid function. We use Bradley-Terry model [2] to capture pairwise preferences between teacher and student response. The proposed generative adversarial training framework allows the student to learn on-policy from its own generated responses via discriminator feedback, eliminating the need to access the teacher’s internal representations.
@@ 2.2 Training
Since the sampling operation in $G(x)$ is non-differentiable with respect to the student model parameters, we treat $D(G(x))$ as a reward and optimize it using policy gradient [33] with established reinforcement learning algorithms. We employ GRPO [27] to train the student in our experiments, with detailed formulations provided in Appendix A.1. For the discriminator $D$, we minimize its training loss derived from Equation (1):
@@ 2.3 Implement GAD with Reinforcement Learning Frameworks
In our experiments, we implement GAD using existing reinforcement learning frameworks, such as verl [28]. GRPO [27] is used as the policy gradient algorithm, which is detailed in Appendix A.1.
@@
Unlike vanilla reinforcement learning, GAD also needs to jointly update the discriminator (i.e., reward model). The discriminator is trained with Bradley-Terry loss on preference pairs to score the teacher response higher than the student’s output, similar to the reward model in RLHF [23]. While conventional RLHF trains a fixed reward model prior to policy optimization which is prone to reward hacking, our approach updates the reward model (discriminator) online to adapt it to the current policy continually.
@@
Term Correspondence Policy Model Generator (i.e.,Student LLM)
@@ 3.1 Setup
Given a dataset of instruction prompts, we collect corresponding responses from a teacher model and use them to distill student models. For the following experiments, we use LMSYS-Chat-1M-Clean222 https://huggingface.co/datasets/OpenLeecher/lmsys_chat_1m_clean, a clean version of the LMSYS-Chat-1M dataset [44]. The dataset is derived from high-quality conversational data collected via the Chatbot Arena333 https://lmarena.ai platform.
@@
We adopt GPT-5-Chat [22] as the teacher model. It is a closed-source chat model ranked ninth on the Chatbot Text Arena leaderboard at the time of writing. For student models, we use the instruction-tuned variants of open-source models from the Qwen2.5 [41] family (Qwen2.5-3B-Instruct, Qwen2.5-7B-Instruct, Qwen2.5-14B-Instruct) and the Llama3 [9] family (Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct).
@@
For training data, we sample 200K samples from LMSYS-Chat-1M-Clean and collect the corresponding GPT-5-Chat responses to the instructions as teacher responses. All models are trained for 3 epochs with a batch size of 256, totaling approximately 2400 optimization steps. The PPO mini-batch size for each policy update is also 256. The maximum context length is set to 2048 tokens for instruction prompts and 1536 for model responses. The training temperature is set to $0.8$. We save checkpoints every 50 steps. More training details can be found in Appendix A.2.
@@
We reserve 500 samples of LMSYS-Chat-1M-Clean as the primary test set. We also include test datasets consisting of a 500-sample subset split from Dolly [6], the 252-sample SelfInst dataset [37], and the 80-question Vicuna benchmark [3] to evaluate out-of-distribution generalization. We report the GPT-4o evaluation scores [45, 10], where GPT-4o first generates reference answers and then scores the output of the student model against them. We also conduct human evaluations on the LMSYS-Chat-1M-Clean test set for qualitative assessment. We select the checkpoint that achieved the highest GPT-4o score and whose response length is within an acceptable range for each experiment. Detailed evaluation protocols are described in Appendix A.3.
@@ 3.2 Main Results
We report the results of automatic evaluation using GPT-4o scores in Figure˜1 and Table˜2. We compare GAD with the instruct model before distillation and the SeqKD baseline. Across all datasets, GAD consistently outperforms the baselines. As shown in Figure˜1, on the LMSYS-Chat test set, Qwen2.5-3B-Instruct trained with GAD matches the performance of Qwen2.5-7B-Instruct trained with SeqKD; similarly, Qwen2.5-7B-Instruct with GAD rivals Qwen2.5-14B-Instruct with SeqKD, and Qwen2.5-14B-Instruct with GAD is comparable to the GPT-5-Chat teacher. In addition, GAD shows particularly strong gains on out-of-distribution generalization benchmarks. On Dolly, SelfInst, and Vicuna, SeqKD yields marginal or even negative improvements, whereas GAD maintains robust performance gains. We attribute this to the superior generalization ability of reinforcement learning compared to supervised fine-tuning [4, 39]. We also provide additional automatic evaluation results in Section˜B.1.
@@ 3.3 Analysis
We evaluate the similarity of local patterns between the student and teacher on the LMSYS-Chat test set in Figure˜5, measured by the F1 score of N-gram overlap. The student is trained from Qwen2.5-14B-Instruct, and the teacher is GPT-5-Chat. The SeqKD student exhibits a higher N-gram overlap while a lower GPT-4o evaluation score compared to the GAD student. This suggests that supervised fine-tuning tends to memorize local lexical patterns [4, 39], whereas our RL-based approach better captures the teacher’s global stylistic characteristics.
@@
We simulate the optimizing patterns of GAD and SeqKD in a toy experiment shown in Figure˜5. We observe that GAD tends to learn reachable modes of the teacher, whereas SeqKD aims to cover all modes. The setup simulates a black-box distillation scenario. We define a discrete Gaussian mixture distribution as a teacher distribution $p$, which has categorical outputs $0,\dots,9$}. A student, modeled as a single Gaussian distribution, learns to imitate the teacher using only output samples without access to $p$. We compare two student training schemes, SeqKD and GAD. The GAD student is optimized using the REINFORCE algorithm [38]. As illustrated in Figure˜5, the SeqKD student exhibits a mode-covering behavior, spreading probability mass across all possible outputs [10]. In contrast, the GAD student focuses on mode-seeking, concentrating probability optimization on reachable regions. We find that such mode-seeking behavior leads to more effective knowledge distillation in LLMs.
@@ 3.4 Ablations
We ablate the choice of discriminator loss in Section˜3.4, and observe that our default Bradley-Terry loss outperforms cross-entropy loss in overall GPT-4o evaluation score. The Bradley-Terry loss is defined in Equation (3), while the cross-entropy loss is a binary classification loss commonly adopted for discriminator in prior works [8, 43, 14, 7]. The cross-entropy discriminator loss can be written as:
@@ 4 Related Work
White-box knowledge distillation of LLM assumes full access to the internal representations or token-level probabilities of a teacher model. Standard white-box approaches align the forward KLD of distribution [17, 30], reverse KLD of distribution [10], hidden states [15, 31] or attention scores [36, 35] between the teacher and the student. Recent work [10, 19, 1] also proves the importance of on-policy distillation where the student learns from its own responses. Such approaches effectively compress large models while preserving semantic similarity. Despite their effectiveness, these methods rely on full teacher access, which is impractical for proprietary LLMs and limits their applicability to closed-source or API-only teachers.
@@
Black-box distillation trains a student model using only the textual outputs of a teacher, typically obtained by API queries to closed-source models such as GPT-5 and Gemini 2.5 [22, 5]. In this setting, conventional white-box distillation methods become infeasible because of the lack of access to the teacher’s logits or hidden representations. The standard approach for this scenario, SeqKD, performs supervised fine-tuning (SFT) on the teacher’s responses [16, 24, 46, 34, 3] to imitate the teacher’s behaviors. Recent work [20, 11, 42, 12] extends this paradigm by performing SFT on the teacher’s reasoning traces to improve the student’s reasoning ability.